OpenStudy (anonymous):

What is the x-intercept in this linear equation? How do I find it? 2y - 4x = -12

OpenStudy (amistre64):

the x axis is the y=0 axis; so the x axis is intercepted (gets crossed) when y=0

OpenStudy (anonymous):

1. re-arrange for y=stuff, solve for y \[2y - 4x = -12 \rightarrow 2y=4x-12 \rightarrow y=2x-6\] 2. when y=0, that is when the output of the function is zero, that's when you'll intercept the x-axis Make sense?

OpenStudy (anonymous):

Yeah, thanks guys.